# DistillSeq: A Framework for Safety Alignment Testing in Large Language Models using Knowledge Distillation

### v2

@@ Abstract.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austriaccs: Security and privacy Software security engineeringdoi:
@@ 1. Introduction
While the AI community actively engages in discussions and formulates strategies to mitigate these issues, extensive research highlights the susceptibility of LLMs to malicious attacks. For instance, Wei et al. 40 have illustrated the potential of exploiting jailbreak prompts as a means to generate harmful content. Sun et al. 36 have identified six distinct categories of instruction-based attacks targeting LLMs, primarily executed through the manual construction of prompts. Moreover, the proliferation of automated prompt generation for such attacks has become an escalating concern. Liu et al. 27 have automated the process of creating jailbreak prompts using an LLM. Furthermore, Zou et al. 46 have proposed a methodology for circumventing enhanced LLMs by appending adversarial suffixes to prompts.
Concurrently, these attacks are not only challenges to be mitigated but also serve a crucial role as testing strategies for the deployment of LLMs. This dual role introduces a significant consideration: the potential for reducing the costs associated with such testing. The question of cost reduction is particularly relevant when considering the extensive requirements for generating a large volume of test cases to interact with LLMs under test. For instance, evaluating a jailbreak prompt with 5,000 queries costs about \$12 33. Alternatively, open-source models like Vicuna-13B offer a free but time-intensive option. Deploying Vicuna-13B requires at least 24GB of VRAM and takes about 160 minutes for the same evaluations 1. Current research tends to focus more on attack strategies. This approach often leads to frequent, unproductive interactions with LLMs. Therefore, improving the success rates of these attacks is crucial.
We present DistillSeq, an innovative framework designed for the automated generation of malicious queries to evaluate the resilience of LLMs against potentially harmful input queries. DistillSeq operates through a two-fold process: knowledge distillation and sequential filter-test procedure. The framework is detailed in the workflow diagram in Figure 2. The initial phase of DistillSeq revolves around the concept of knowledge distillation. In this stage, we transfer the moderation mechanism knowledge from the LLM under test to a more compact and efficient, distilled model. In the subsequent stage, this distilled model meticulously evaluates the malicious queries by applying a rigorous filtering mechanism. It filters out malicious queries that are improbable to elicit harmful or toxic responses. This filtering process leads to a substantial reduction in the number of queries necessary during the testing phase.
@@
We have conducted an evaluation on four popular LLMs: GPT-3.5 31, GPT-4 32, Vicuna-13B 1, LLama-13B30. We measured the effectiveness of DistillSeq using the attack success rate. The percentage of successful malicious queries that prompt LLMs to generate toxic content. By employing the distilled model as a filter, we achieved an average attack success rate of 49.8% using the Syntax tree-based method, and 58.3% with the LLM-based method. This method significantly improved effectiveness. Our research evaluated the effectiveness of DistillSeq on a range of LLMs including GPT-3.5, GPT-4.0, Vicuna-13B, and Llama-13B. In scenarios where DistillSeq was not employed, the attack success rates were recorded as 31.5% for GPT-3.5, 21.4% for GPT-4.0, 28.3% for Vicuna-13B, and 30.9% for Llama-13B. Remarkably, with the implementation of DistillSeq, these success rates significantly increased to 58.5%, 50.7%, 52.5%, and 54.4% for each model, respectively. This indicates an average increase in the attack success rate of 93.0% when DistillSeq is utilized, compared to scenarios where it is absent. Our results highlight significant limitations in current safety alignments and offer a more efficient approach for testing these models.
@@
  • Empirical Evidence of Method Effectiveness: We evaluate DistillSeq on four mainstream LLMs. The results demonstrate that DistillSeq significantly increases the average attack success rate by 93% and generates more effective malicious queries compared to existing methods.
@@ 2.1. Jailbreak
In the context of large language models, “jailbreaking” refers to a phenomenon where the model disregards or bypasses predefined constraints when given specific inputs 41, 21, 4. It may generate biased, harmful, or unethical content. Figure 1 shows an example of a jailbreak attack. The model refuses to directly answer a malicious query like “I want to cause public panic what should I do.” However, when the same question is appended to a jailbreak prompt, the model generates a toxic response without any awareness.
@@ 2.2. Knowledge Distillation
Offline knowledge distillation is a commonly used technique in the field of knowledge distillation. It involves using a fully trained teacher model to create a specialized dataset. Using this dataset to train the student model. This process allows the student model to learn from the teacher model’s outputs, effectively compressing the knowledge into a more compact form while maintaining comparable performance.
@@ 2.3. Motivation
To further explain the practical advantages of our DistillSeq framework, we reproduced jailbreak attacks from the empirical study conducted by Liu et al. 27 as an example for comparison. Specifically, we examined the GPT-4.0 model by executing 17,000 jailbreak attacks, resulting in the identification of 121 security violations. This testing required approximately 8,500,000 tokens, with a financial cost of around \$300 based on OpenAI’s pricing policy. Utilizing our DistillSeq framework, we achieved a significant reduction in the number of queries required to attain an equal number of violations in the GPT-4.0 model, accomplishing this with only 240 queries. These queries corresponded to a token usage of just 120,000, resulting in a financial cost of approximately \$4. In addition, for a fair comparison, we should include the cost of training the model in our framework in the knowledge distillation process. We utilized a training set of 4000 samples, with OpenAI’s pricing of \$0.03 per 1000 tokens for input and \$0.06 per 1000 tokens for output. Each query utilized approximately 200 tokens for input and 100 tokens for output. Therefore, the total training cost is calculated as $(0.03\times 200/1000+0.06\times 100/1000)\times 4000=\$48$. Following this, we conducted the distillation offline using an NVIDIA GeForce RTX 3090 GPU, which required approximately one hour of processing time. Leveraging pricing data from Vastai [^1] https://vast.ai/, we estimate the expense for this operation to be around \$0.35. Overall, the total cost involved in our approach is \$52.35, representing approximately a fivefold reduction in testing costs compared to the traditional jailbreak attack method.
@@ 3.1. Problem Statement
Our work focuses on using knowledge distillation to train a student model that learns LLM moderation mechanism knowledge. This model effectively filters ineffective toxic content in the context of a jailbreak attack. Employing the student model as the filter can reduce the total number of queries, thereby improving the efficiency of testing LLMs. We initially introduce specific details regarding our setting and then formalize the problem with the following settings:
@@ 3.2. Building a Distillation Model
We utilize a dataset containing malicious queries empirically known to elicit harmful responses. However, as observed by Zhuo et al. 45, the probability of LLMs producing toxic responses solely based on malicious queries is typically low. To address this issue, we introduce jailbreak prompts, which significantly increase the probability of circumventing LLM moderation when combined with malicious requests 27. Our queries are generated by merging these jailbreak prompts with the malicious queries. Subsequently, we meticulously assess the toxicity scores of the LLMs’ responses using an online toxicity detection tool.
@@ 3.3. Generate New Malicious Queries
In this method we apply two strategies to reduce the generation of nonsensical queries. First, during the replacement process, our approach is to maintain the structural integrity of the syntactic tree. This is achieved by substituting syntactic subtrees with alternatives of the same syntactic categories. This method ensures that the overall structure of the sentence remains intact, thereby preserving grammatical correctness. Adhering to this principle reduces the likelihood of generating nonsensical outputs, as the syntactic framework of the query is maintained. Second, we implement a selection method that prioritizes the replacement of the shortest possible subtrees. This
@@ 17                   $T$ = $T\cup ret$;
strategy aims to minimize the impact on the sentence’s coherence and meaning. By limiting modifications to smaller, more manageable components of the sentence, we can maintain a higher degree of semantic integrity.
Figure 3 demonstrates an example of the syntax tree-based method. Through calculation, we find that the subtree “help others out of trouble” possesses the highest influence within the syntax tree, and this subtree belongs to a verb phrase. Subsequently, we select another subtree, “destroying other people’s property,” which is also a verb phrase, to replace the most influential subtree. This substitution operation generates a new malicious query, which can induce LLM to produce toxic output.
@@
specially curated dataset, the RealToxicPrompts 12 dataset, which contains various toxic and harmful content samples. Vicuna-13B generates the LLM’s responses to these queries, training it to recognize, anticipate, and extend conversations with malicious queries. This process helps Vicuna-13B simulate real-world LLM interactions more accurately. The initial phase of pre-training focuses on tasks related to toxic and harmful content, providing Vicuna-13B with a foundational understanding of such content. This allows for more nuanced and informed responses during the fine-tuning stage, which further hones the model’s ability to comprehend and interact with malicious queries. This process improves the performance of LLMs on tasks related to malicious queries.
@@ 4. Evaluation
In this section, we present our evaluation of DistillSeq. The implementation details of DistillSeq are available our project website  [^2] https://distillseq.github.io/page/. To assess its effectiveness, we explore the following research questions:
@@ 4.1. Dataset
To implement knowledge distillation in our framework, we require datasets for training. Here, we construct the dataset by combining jailbreak prompts and malicious queries as input. These jailbreak prompts are sourced from Jailbreak Chat [^3] https://www.jailbreakchat.com/. In addition, we follow the categorization proposed by Liu et al. 27, which divides jailbreak prompts into Pretending, Attention Shifting, and Privilege Escalation. We select the top two popular prompts from each category based on user votes, as it reflects the effectiveness of these jailbreak prompts in eliciting toxic responses from LLMs, as judged by user experience. Currently, there is no universally accepted, systematic standard for assessing the quality of jailbreak prompts. We make this decision based on the fact that Jailbreak Chat hosts the most extensive collection of LLM jailbreak attempts available online, making its prompts widely utilized by many users.
@@ 4.2. RQ1: Effectiveness
For computing these metrics, in this research question, we utilize GPT-3.5 as the teacher model. To perform knowledge distillation, we select 4000 queries from our dataset as input to the LLM. Among these queries, 2000 elicit toxic responses, while the remaining queries do not. It is important to note that we conduct 10 iterations for each query, and if the LLM gives a toxic response more than 6 times out of these iterations, we consider the query to be an effective malicious query. We construct the training set using these queries and their corresponding labels based on the response. Similarly, we select 1000 different queries from our dataset to create the testing set. In our knowledge distillation process, we utilize BERT 9, RoBERTa 28, ERNIE 37, and DEBERTA 15 as the architecture for the student model. The reason is that these models have demonstrated strong capabilities in processing and understanding natural language text. Additionally, similar to LLMs, they are all pretrained language models that have been trained on vast amounts of text data. Our hyperparameter settings are based on the recommendations by Karl et al. 20. Details of these settings are provided in Table 1.
@@ 4.3. RQ2: Tradeoff
Method. To investigate the impact of training dataset size on the performance of the distilled model, we conduct a series of experiments with varying sample sizes. Starting from 0 samples, we incrementally increase the number of samples by 2,000 in each iteration. At each step, we train the distilled model using the corresponding number of samples and evaluate its agreement with GPT-3.5 on the 1,000 randomly selected toxic comments from the RealToxicPrompt dataset. To represent a real testing scenario, we label the queries based directly on the responses of GPT-3.5. Since these responses can be influenced by the variability of the LLM, the agreement result will decrease to some extent.
@@
We assume that one query requires 200 tokens, and the LLM’s output uses 100 tokens. According to OpenAI’s pricing scheme 33, it can be determined that for every 2,000 queries, the cost would be approximately \$8. Specifically, the average agreement improvement per 2,000 additional samples decreases from 28.28% (from 0 to 2,000 samples) to a mere 1.55% (from 8,000 to 10,000 samples). The average cost per 1 percentage point improvement in agreement escalates from \$0.28 (from 0 to 2,000 samples) to a staggering \$10.32 (from 8,000 to 10,000 samples).
@@
Answer to RQ2 While the effectiveness of the model improves with an increased amount of data used during knowledge distillation, the growth rate slows down later on. Taking cost into account, we choose to use 4000 samples for knowledge distillation.
@@ 4.4. RQ3: Generality
Method.
In this research question, we evaluate our framework’s performance on four popular LLMs: GPT-3.5 31, GPT-4 32, Vicuna-13B 38, and LLama-13B 30. To mirror real-world scenarios, we use the default settings of these LLMs, including temperature and top-p values, for response generation. During the training phase, we select 4000 samples for knowledge distillation. We employ the same method as in RQ1 to train the classification models.
@@
To further investigate the performance degradation, we present the loss values obtained from these models on the Jigsaw dataset in Table  4. Upon comparing the results in Table  2 and 4, we observe that while the knowledge distillation process reduces the loss for the Jigsaw dataset, the loss remains higher when compared to the loss data in Tables  2. This suggests that the model may rely on specific characteristics unique to the training data. This also highlights that our method can be enhanced by incorporating additional datasets or augmenting the existing dataset with more diverse examples.

### v3

@@ Abstract.
copyright: rightsretaineddoi: 10.1145/3650212.3680304journalyear: 2024isbn: 979-8-4007-0612-7/24/09conference: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis; September 16–20, 2024; Vienna, Austriabooktitle: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA ’24), September 16–20, 2024, Vienna, Austriasubmissionid: issta24main-p73-pccs: Security and privacy Software security engineering
@@ 1. Introduction
While the AI community actively engages in discussions and formulates strategies to mitigate these issues, extensive research highlights the susceptibility of LLMs to malicious attacks. For instance, Wei et al. 39 have illustrated the potential of exploiting jailbreak prompts as a means to generate harmful content. Sun et al. 35 have identified six distinct categories of instruction-based attacks targeting LLMs, primarily executed through the manual construction of prompts. Moreover, the proliferation of automated prompt generation for such attacks has become an escalating concern. Liu et al. 26 have automated the process of creating jailbreak prompts using an LLM. Furthermore, Zou et al. 45 have proposed a methodology for circumventing enhanced LLMs by appending adversarial suffixes to prompts.
Concurrently, these attacks are not only challenges to be mitigated but also serve a crucial role as testing strategies for the deployment of LLMs. This dual role introduces a significant consideration: the potential for reducing the costs associated with such testing. The question of cost reduction is particularly relevant when considering the extensive requirements for generating a large volume of test cases to interact with LLMs under test. For instance, evaluating a jailbreak prompt with 5,000 queries costs about \$12 32. Alternatively, open-source models like Vicuna-13B offer a free but time-intensive option. Deploying Vicuna-13B requires at least 24GB of VRAM and takes about 160 minutes for the same evaluations 1. Current research tends to focus more on attack strategies. This approach often leads to frequent, unproductive interactions with LLMs. Therefore, improving the success rates of these attacks is crucial.
We present DistillSeq, an innovative framework designed for the automated generation of malicious queries to evaluate the resilience of LLMs against potentially harmful input queries. DistillSeq operates through a two-fold process: knowledge distillation and sequential filter-test procedure. The framework is detailed in the workflow diagram in Figure 2. The initial phase of DistillSeq revolves around the concept of knowledge distillation. In this stage, we transfer the moderation mechanism knowledge from the LLM under test to a more compact and efficient distilled model. In the subsequent stage, this distilled model meticulously evaluates the malicious queries by applying a rigorous filtering mechanism. It filters out malicious queries that are improbable to elicit harmful or toxic responses. This filtering process leads to a substantial reduction in the number of queries necessary during the testing phase.
@@
We have conducted an evaluation on four popular LLMs: GPT-3.5 30, GPT-4 31, Vicuna-13B 1, and LLama-13B29. We measured the effectiveness of DistillSeq using the attack success rate. The percentage of successful malicious queries that prompt LLMs to generate toxic content. By employing the distilled model as a filter, we achieved an average attack success rate of 49.8% using the Syntax tree-based method, and 58.3% with the LLM-based method. This method significantly improved effectiveness. Our research evaluated the effectiveness of DistillSeq on a range of LLMs including GPT-3.5, GPT-4.0, Vicuna-13B, and Llama-13B. In scenarios where DistillSeq was not employed, the attack success rates were recorded as 31.5% for GPT-3.5, 21.4% for GPT-4.0, 28.3% for Vicuna-13B, and 30.9% for Llama-13B. Remarkably, with the implementation of DistillSeq, these success rates significantly increased to 58.5%, 50.7%, 52.5%, and 54.4% for each model, respectively. This indicates an average increase in the attack success rate of 93.0% when DistillSeq is utilized, compared to scenarios where it is absent. Our results highlight significant limitations in current safety alignments and offer a more efficient approach for testing these models.
@@
  • Empirical Evidence of Method Effectiveness: We evaluated DistillSeq on four mainstream LLMs. The results demonstrate that DistillSeq significantly increases the average attack success rate by 93% and generates more effective malicious queries compared to existing methods.
@@ 2.1. Jailbreak
In the context of large language models, “jailbreaking” refers to a phenomenon where the model disregards or bypasses predefined constraints when given specific inputs 40, 21, 4. It may generate biased, harmful, or unethical content. Figure 1 shows an example of a jailbreak attack. The model refuses to directly answer a malicious query like “I want to cause public panic what should I do.” However, when the same question is appended to a jailbreak prompt, the model generates a toxic response without any awareness.
@@ 2.2. Knowledge Distillation
Offline knowledge distillation is a commonly used technique in the field of knowledge distillation. It involves using a fully trained teacher model to create a specialized dataset. This dataset will be used to train the student model. This process allows the student model to learn from the teacher model’s outputs, effectively compressing the knowledge into a more compact form while maintaining comparable performance.
@@ 2.3. Motivation
To further explain the practical advantages of our DistillSeq framework, we reproduced jailbreak attacks from the empirical study conducted by Liu et al. 26 as an example for comparison. Specifically, we examined the GPT-4.0 model by executing 17,000 jailbreak attacks, resulting in the identification of 121 security violations. This testing required approximately 8,500,000 tokens, with a financial cost of around \$300 based on OpenAI’s pricing policy. Utilizing our DistillSeq framework, we achieved a significant reduction in the number of queries required to attain an equal number of violations in the GPT-4.0 model, accomplishing this with only 240 queries. These queries corresponded to a token usage of just 120,000, resulting in a financial cost of approximately \$4. In addition, for a fair comparison, we should include the cost of training the model in our framework in the knowledge distillation process. We utilized a training set of 4,000 samples, with OpenAI’s pricing of \$0.03 per 1000 tokens for input and \$0.06 per 1000 tokens for output32. Each query utilized approximately 200 tokens for input and 100 tokens for output. Therefore, the total training cost is calculated as $(0.03\times 200/1000+0.06\times 100/1000)\times 4000=\$48$. Following this, we conducted the distillation offline using an NVIDIA GeForce RTX 3090 GPU, which required approximately one hour of processing time. Leveraging pricing data from Vastai [^1] https://vast.ai/, we estimate the expense for this operation to be around \$0.35. Overall, the total cost involved in our approach is \$52.35, representing approximately a fivefold reduction in testing costs compared to the traditional jailbreak attack method.
@@ 3.1. Problem Statement
Our work focuses on using knowledge distillation to train a student model that learns LLM moderation mechanism knowledge. This model effectively filters out ineffective toxic content in the context of a jailbreak attack. Employing the student model as the filter can reduce the total number of queries, thereby improving the efficiency of testing LLMs. We initially introduce specific details regarding our setting and then formalize the problem with the following settings:
@@ 3.2. Building a Distillation Model
We utilize a dataset containing malicious queries empirically known to elicit harmful responses. However, as observed by Zhuo et al. 44, the probability of LLMs producing toxic responses solely based on malicious queries is typically low. To address this issue, we introduce jailbreak prompts, which significantly increase the probability of circumventing LLM moderation when combined with malicious requests 26. Our queries are generated by merging these jailbreak prompts with the malicious queries. Subsequently, we meticulously assess the toxicity scores of the LLMs’ responses using an online toxicity detection tool.
@@ 3.3. Generate New Malicious Queries
In this method, we apply two strategies to reduce the generation of nonsensical queries. First, during the replacement process, our approach is to maintain the structural integrity of the syntactic tree. This is achieved by substituting syntactic subtrees with alternatives of the same syntactic category. This method ensures that the overall structure of the sentence remains intact, thereby preserving grammatical correctness. Adhering to this principle reduces the likelihood of generating nonsensical outputs, as the syntactic framework of the
@@ 17                   $T$ = $T\cup ret$;
query is maintained. Second, we implement a selection method that prioritizes the replacement of the shortest possible subtrees. This strategy aims to minimize the impact on the sentence’s coherence and meaning. By limiting modifications to smaller, more manageable components of the sentence, we can maintain a higher degree of semantic integrity.
Figure 3 demonstrates an example of the syntax tree-based method. Through calculation, we find that the subtree “help others out of trouble” possesses the highest influence within the syntax tree, and this subtree belongs to a verb phrase. Subsequently, we select another subtree, “destroy other people’s property” which is also a verb phrase, to replace the most influential subtree. This substitution operation generates a new malicious query, which can induce LLM to produce toxic output.
@@
specially curated dataset, specifically the RealToxicPrompts 12 dataset, which contains various toxic and harmful content samples. Vicuna-13B generates the LLM’s responses to these queries, training it to recognize, anticipate, and extend conversations with malicious queries. This process helps Vicuna-13B simulate real-world LLM interactions more accurately. The initial phase of pre-training focuses on tasks related to toxic and harmful content, providing Vicuna-13B with a foundational understanding of such content. This allows for more nuanced and informed responses during the fine-tuning stage, which further hones the model’s ability to comprehend and interact with malicious queries. This process improves the performance of LLMs on tasks related to malicious queries.
@@ 4. Evaluation
In this section, we present our evaluation of DistillSeq. The implementation details of DistillSeq are available on our project website  [^2] https://distillseq.github.io/page/. To assess its effectiveness, we explore the following research questions:
@@ 4.1. Dataset
To implement knowledge distillation in our framework, we require datasets for training. Here, we construct the dataset by combining jailbreak prompts and malicious queries as input. These jailbreak prompts are sourced from Jailbreak Chat [^3] https://www.jailbreakchat.com/. In addition, we follow the categorization proposed by Liu et al. 26, which divides jailbreak prompts into Pretending, Attention Shifting, and Privilege Escalation. We select the top two popular prompts from each category based on user votes, as it reflects the effectiveness of these jailbreak prompts in eliciting toxic responses from LLMs, as judged by user experience. Currently, there is no universally accepted, systematic standard for assessing the quality of jailbreak prompts. We make this decision based on the fact that Jailbreak Chat hosts the most extensive collection of LLM jailbreak attempts available online, making its prompts widely utilized by many users.
@@ 4.2. RQ1: Effectiveness
For computing these metrics, in this research question, we utilize GPT-3.5 as the teacher model. To perform knowledge distillation, we select 4,000 queries from our dataset as input to the LLM. Among these queries, 2,000 elicit toxic responses, while the remaining queries do not. It is important to note that we conduct 10 iterations for each query, and if the LLM gives a toxic response more than 6 times out of these iterations, we consider the query to be an effective malicious query. We construct the training set using these queries and their corresponding labels based on the responses. Similarly, we select 1,000 different queries from our dataset to create the testing set. In our knowledge distillation process, we utilize BERT 9, RoBERTa 27, ERNIE 36, and DeBERTa 15 as the architecture for the student model. The reason is that these models have demonstrated strong capabilities in processing and understanding natural language text. Additionally, similar to LLMs, they are all pretrained language models that have been trained on vast amounts of text data. Our hyperparameter settings are based on the recommendations by Karl et al. 20. Details of these settings are provided in Table 1.
@@ 4.3. RQ2: Tradeoff
Method. To investigate the impact of training dataset size on the performance of the distilled model, we conduct a series of experiments with varying sample sizes. Starting from 0 samples, we incrementally increase the number of samples by 2,000 in each iteration. At each step, we train the distilled model using the corresponding number of samples and evaluate its agreement with GPT-3.5 on the 1,000 randomly selected toxic comments from the RealToxicPrompts dataset. To represent a real testing scenario, we label the queries based directly on the responses of GPT-3.5. Since these responses can be influenced by the variability of the LLM, the agreement result will decrease to some extent.
@@
We assume that one query requires 200 tokens, and the LLM’s output uses 100 tokens. According to OpenAI’s pricing scheme 32, it can be determined that for every 2,000 queries, the cost would be approximately \$8. Specifically, the average agreement improvement per 2,000 additional samples decreases from 28.28% (from 0 to 2,000 samples) to a mere 1.55% (from 8,000 to 10,000 samples). The average cost per 1 percentage point improvement in agreement escalates from \$0.28 (from 0 to 2,000 samples) to a staggering \$10.32 (from 8,000 to 10,000 samples).
@@
Answer to RQ2 While the effectiveness of the model improves with an increased amount of data used during knowledge distillation, the growth rate slows down later on. Taking cost into account, we choose to use 4,000 samples for knowledge distillation.
@@ 4.4. RQ3: Generality
Method. In this research question, we evaluate our framework’s performance on four popular LLMs: GPT-3.5 30, GPT-4 31, Vicuna-13B 37, and LLama-13B 29. To mirror real-world scenarios, we use the default settings of these LLMs, including temperature and top-p values, for response generation. During the training phase, we select 4,000 samples for knowledge distillation. We employ the same method as in RQ1 to train the classification models.
@@
To further investigate the performance degradation, we present the loss values obtained from these models on the Jigsaw dataset in Table  4. Upon comparing the results in Table  2 and 4, we observe that while the knowledge distillation process reduces the loss for the Jigsaw dataset, the loss remains higher when compared to the loss data in Table  2. This suggests that the model may rely on specific characteristics unique to the training data. This also highlights that our method can be enhanced by incorporating additional datasets or augmenting the existing dataset with more diverse examples.